# Single-photon scattering and bound states in a one-dimensional waveguide with topological giant atom

### v1

@@ V REMARKS and CONCLUSIONS
In the current experimental availability, the proposed model can be realized in the superconducting circuits.In 2014, the first experimental realized the coupling between a superconducting transmon quantum qubit and the surface acoustic waves (SAWs) waveguide was realized. Here, the qubit is $20$ times larger than the wavelength of the SAWs in size and it therefore reached the giant atom regime [70]. This ratio between the size of the giant atom and the wavelength was then developed to $100$ and the light-matter interaction was achieved by tens of MHz [34]. Additionally, the giant atom was also achieved by coupling artificial atoms created with Josephson junctions to superconducting circuits through the capacitance or inductance. In Ref. [37], the giant atom was experimentally achieved by and couple it to a waveguide at multiple, yet well separated, discrete locations. The distance between two coupling points could reach $20.54$ mm [44]. Recently, the topology of two nested giant spin ensembles(GSE) has been experimentally demonstrated, the distance between the two coupling points of the inner(outer) is designed to be $8.3\,(16.6)$cm [71]. It provides a new platform for “giant atom” physics. Moreover,the CRW constructed with $9$ superconducting qubits is also realized in the experiment, where the nearest-neighbor coupling strength is $\xi/2\pi=50$ MHz [72]. In a recent experiment, Painter’s group expanded the CRW consist of a $42$ unit-cell array of capacitively coupled lumped-element microwave resonators [73], and the SSH topological structure was also constructed with the hopping strength $J(1\pm\delta)$ where $J/2\pi=368$ MHz, $\delta=0.282$. Furthermore, the qubit-resonator coupling strength in such structure was achieved by $g/2\pi=124.6$ MHz [69]
The establishment of TGA in our work provides an unconventional approach to regulate the transmission and distribution of the single photon in the CRW. Here, the interference effect for the propagation of the photon in the regime of the giant atom will be applicable in quantum control and quantum information processing. We hope that our work of controlling photons through the giant atom systems combined with topologies will stimulate new research of the hybrid system and broaden the range of application of artificial giant atom.

### v2

@@ V REMARKS and CONCLUSIONS
In the current experimental availability, the proposed model can be realized in the superconducting circuits. In 2014, the first experiment which couples the superconducting transmon quantum qubit and the surface acoustic waves (SAWs) waveguide was realized. Here, the qubit is $20$ times larger than the wavelength of the SAWs in size and it therefore reached the giant atom regime gus2014 . This ratio between the size of the giant atom and the wavelength was then developed to $100$ and the light-matter interaction was achieved by tens of MHz andg2019 . Additionally, the giant atom was also achieved by coupling artificial atoms created with Josephson junctions to superconducting circuits through the capacitance or inductance. In Ref. Kanb2020 , the giant atom was experimentally coupled to a waveguide at multiple, yet well separated locations. The distance between two coupling points could reach $20.54$ mm vada2021 . Recently, the topology of two nested giant spin ensembles (GSE) has been experimentally demonstrated, the distance between the two inner (outer) coupling points of the is designed to be $8.3\,(16.6)$cm wangzq2022 . It provides a new platform for “giant atom” physics. Moreover,the CRW constructed with $9$ superconducting qubits is also realized in the experiment, where the nearest-neighbor coupling strength is $\xi/2\pi=50$ MHz rou2017 . In a recent experiment, Painter’s group expanded the CRW consist of a $42$ unit-cell array of capacitively coupled lumped-element microwave resonators zhangxy2023 , and the SSH topological structure was also constructed with the hopping strength $J(1\pm\delta)$ where $J/2\pi=368$ MHz, $\delta=0.282$. Furthermore, the qubit-resonator coupling strength in such structure was achieved by $g/2\pi=124.6$ MHz kim2021
The establishment of TGA in our work provides an unconventional approach to regulate the transmission and distribution of the single photon in the CRW. Here, the interference effect for the propagation of the photon in the regime of the giant atom will be applicable in quantum control and quantum information processing. We hope that our work of controlling photons through the giant atom systems combined with topologies will stimulate further research of the hybrid system and broaden the range of application of artificial giant atom.